An object from rest starts to accelerate at 2.5m/s-2


Determine the time required to cover 100m.


1
Expert's answer
2021-03-21T11:37:18-0400

"s=\frac{at^2}{2}\implies t=\frac{2s}{a},"

"t=\frac{2\cdot 100}{2.5}=80~s."